# Calculating the EFT likelihood via saddle-point expansion

### v1

@@ 4 Example: the EFT likelihood with Gaussian distribution
  • It has been demonstrated that the saddle-point expansion method in false vacuum decay theory can be seamlessly applied to the path-integral approach of LSS after some appropriate modifications and re-derivations. Although we have met the negative modes, and even complex modes of $S^{\prime\prime}_{g}[\bar{\boldsymbol{\phi}}_{g}]$, because all the eigenvalues of the gradient flow eigenequations Eq. (4.14) are real and positive ($m_{n}\in\mathbb{R},m_{n}>0$), the result is ensured to be real through the path integral, consistent with the argument as presented in Sec. 2.1. This aspect of the result is very crucial, as it directly relates to its physical interpretation as a probability.
  • Although we have examined only one specific model, the rationale for applying the gradient flow equation to determine the steepest descent contour, as we discussed below, is general and can be applied to any kind of more complicated models. After this, we can compute the arbitrary likelihood of the form $\int\mathcal{D}\phi\,{\rm e}^{-S[\phi]}$ up to the second-order near the saddle points. This approach could be instrumental in providing more accurate predictions for the large-scale structure of the universe.
@@
$\wp[\delta]$ $={\rm ln}\,\mathcal{P}[\delta]=-{\rm ln}\,\left(\frac{\sqrt{\pi}D_{1}}{2}\right)S[\bar{\boldsymbol{\phi}}]\,,$ (4.19)
@@
Since neither the matrices $\mathcal{M}$ nor $\mathcal{M}_{g}$ depend on the overdensity fields $\delta$ and $\delta_{g}$, the eigenvalues will be field-independent. We can also see this in the expression of $m_{n}$ in Appendix B, where we provide a set of approximate solutions for them. Thus the logarithmic terms in Eq. (4.19) and Eq.(4.20) will be absorbed into the normalization factors. Consequently, when defining the conditional likelihood of them, the terms appearing in the conditional likelihood will only include the saddle-point contribution of the two actions, which is
@@ 5 Summary and conclusion
The path-integral approach of the EFTofLSS offers a novel method for encapsulating cosmological effects into a compact analytical framework, enabling the prediction of cosmological observables. In this work, we aim to apply the saddle-point expansion technique, traditionally employed in quantum field theory for calculating the decay rates, to this context to improve the precision of EFT likelihood calculations. As a result, after appropriately reformulating the likelihood expressions, this task reduces to computing the path integral over the field $\boldsymbol{\phi}$ in three-dimensional Euclidean space. Consequently, by applying the same reasoning as in quantum field theory, we can calculate the likelihood, with the $\phi$ space replaced by the $\boldsymbol{\phi}$ space.
@@
A key challenge in the saddle-point expansion method is the treatment of the negative eigenvalues of the second derivative of the action at the saddle point. In this work, we present a systematic discussion of this aspect in three distinct cases. As an illustrative example, we compute the matter and joint likelihood under the assumption of Gaussian distribution, retaining terms up to $J^{2}$ and $J_{g}^{2}$ order. For the matter likelihood, since there are no negative eigenvalues associated with $S^{\prime\prime}[\boldsymbol{\phi}]$, we can directly apply the Gaussian integral formula to obtain the result. However, when calculating the joint likelihood, the eigenvalues of $S^{\prime\prime}[\boldsymbol{\phi}_{g}]$ are often non-positive and even complex, necessitating an approximation of the original integral contour with a sum of the steepest descent contours in the field space. The treatment of these complex modes goes beyond the scope of the false vacuum decay theory discussion, primarily because the second derivative of the action in the path-integral approach of LSS often takes a specific form, resulting in a general non-Hermitian three-dimensional matrix. Even so, we still demonstrate in Sec. 4 that for this case, the Picard-Lefschetz theory remains applicable in this context by complexifying the action and utilizing the gradient flow equation in the field space. We emphasize here again that although the integrals along each contour may yield complex results, their collective interference leads to a real final result.
@@
Another point is that, as shown in Eq. (4.8) and Eq. (4.18), the result of the saddle-point expansion is essentially equivalent to the saddle-point approximation, differing only by a multiplicative factor. In the example provided in Sec. 4, this factor is independent of $\delta_{g}$ and $\delta$, therefore it will be absorbed into the normalization constant and resulting in the same outcome as in [19], as illustrated in Eq. (4.21). However, this does not imply that the result of the saddle-point expansion is trivial, but rather that the action we have chosen in this example is truncated at $\boldsymbol{\phi}^{2}$ order. Furthermore, the method we present in this paper is general, and can be used to cope with more complicated situations.
@@
Several unresolved issues remain in this work. For instance, we have not obtained an exact analytic solution of Eq.(4.16). The reason is that although the combination of the matrix $\mathcal{M}_{g}$ with its complex conjugate $\mathcal{M}_{g}^{*}$ is Hermitian, as illustrated in Eq. (4.16), solving for this eigenequation is equivalent to solving a cubic equation, which does not have a simple analytic solution. Another important aspect concerns the treatment of higher-order terms. In the theory of false vacuum decay, the Lagrangian is typically truncated at the $\phi^{4}$ order due to the non-renormalizable nature of higher-order terms, which are strongly suppressed by the energy scale of new physics, $\mu$. This raises the question of whether a similar suppression applies to higher-order terms in our case. In the literature [47] and [45], the authors computed the renormalization group equations for the partition function up to $J^{2}$ order and arbitrary $J^{n}$ order ($n\rightarrow\infty$). They found that, given the same initial conditions, the running of the bias parameter $b$ exhibits similar behavior. Does this imply when calculating the EFT likelihood, we can neglect the terms higher than $J^{2}$ (analogous to the non-renormalizable terms in quantum field theory)? In the literature [45] the authors perform a dimension analysis on the terms of all orders of $J$, however, do not draw an analogy to the case of quantum field theory. Then for the higher-order terms of $J$, are they also suppressed by the hard momentum cutoff $\Lambda$ introduced during the regularization process? e.g. consider the term $\int_{\boldsymbol{x}}J^{2}(\boldsymbol{x})O[\delta_{\Lambda}(\boldsymbol{x})]$, should we need to express this as $\frac{C_{*}}{\Lambda^{3+d_{O}}}\int_{\boldsymbol{x}}J^{2}(\boldsymbol{x})O[\delta_{\Lambda}(\boldsymbol{x})]$ such that this term will be suppressed by $\Lambda$, where $d_{O}$ is the dimension of the arbitrary operator $O$ and $C_{*}$ is the dimensionless “coupling constant”? If similar reasoning as in quantum field theory can be applied, then these higher-order terms would contribute little to the partition function. Unfortunately, we have not found relevant discussions in any literature, so we refrain from drawing a definitive conclusion at this stage. Additionally, our calculation assumes a Gaussian initial condition. If the primordial probability distribution deviates from the Gaussian form, then the effective action will include an additional interaction term [46]. For the standard approach [52, 53], this term introduces a cubic interaction term in the field $\boldsymbol{\phi}$ when calculating the likelihood. The saddle-point expansion method would then require a careful treatment of this term. We leave a more detailed investigation of this issue to future work.
@@ Acknowledgments
JYK thanks Ying-Zhuo Li for his useful discussions. We acknowledge the support by the National Science Foundation of China (No. 12147217, 12347163), the China Postdoctoral Science Foundation (No. 2024M761110), and the Natural Science Foundation of Jilin Province, China (No. 20180101228JC).
Appendix A The contribution of the saddle-point
@@
$\frac{\delta S[\boldsymbol{\phi}]}{\delta\boldsymbol{\phi}}\bigg{|}_{\boldsymbol{\phi}=\bar{\boldsymbol{\phi}}}$ $=\mathcal{J}^{a}+\mathcal{M}^{ab}\bar{\boldsymbol{\phi}}^{b}=0\,,$ (A.1)
$\frac{\delta S_{g}[\boldsymbol{\phi}_{g}]}{\delta\boldsymbol{\phi}_{g}}\bigg{|}_{\boldsymbol{\phi}_{g}=\bar{\boldsymbol{\phi}}_{g}}$ $=\mathcal{J}^{a}_{g}+\mathcal{M}^{ab}_{g}\bar{\boldsymbol{\phi}}^{b}_{g}=0\,.$ (A.2)
These two equations can be readily solved once the forms of $\mathcal{M}$ and $\mathcal{M}_{g}$ are known, i.e. through Eq. (4.3) and Eq. (4.4). The solution for the former case is
$\bar{\boldsymbol{\phi}}^{a}$ $=-(\mathcal{M}^{ab})^{-1}\mathcal{J}^{b}$
$=-\begin{pmatrix}\frac{P_{\rm in}^{-1}}{D_{1}^{2}}&\frac{1}{iD_{1}}\\ \frac{1}{iD_{1}}&0\end{pmatrix}\begin{pmatrix}i\delta\\ 0\end{pmatrix}$
$=\begin{pmatrix}\frac{{i\delta}}{D_{1}^{2}P_{\rm in}}\\ \frac{\delta}{D_{1}}\end{pmatrix}.$ (A.3)
@@
$\bar{\boldsymbol{\phi}}_{g}^{a}$ $=-(\mathcal{M}^{ab}_{g})^{-1}\mathcal{J}^{b}_{g}$
$=\frac{1}{-2b_{1}D_{1}^{2}P_{\epsilon_{g}\epsilon_{m}}+P_{\epsilon_{g}}D_{1}^{2}-P_{\epsilon_{g}\epsilon_{m}}P_{\rm in}^{-1}}\times$
@@
$\approx\frac{1+2b_{1}P_{\epsilon_{g}\epsilon_{m}}/P_{\epsilon_{g}}}{D_{1}^{2}P_{\epsilon_{g}}}\begin{pmatrix}iD_{1}^{2}\delta_{g}-ib_{1}D_{1}^{2}\delta-iP_{\epsilon_{g}\epsilon_{m}}P_{\rm in}^{-1}\delta\\ i\delta_{g}(-b_{1}D_{1}^{2}-P_{\epsilon_{g}\epsilon_{m}}P_{\rm in}^{-1})+i\delta(D_{1}^{2}b_{1}^{2}+P_{\epsilon_{g}}P_{\rm in}^{-1})\\ -D_{1}P_{\epsilon_{g}\epsilon_{m}}\delta_{g}+i\delta(ib_{1}D_{1}P_{\epsilon_{g}\epsilon_{m}}-iD_{1}P_{\epsilon_{g}})\end{pmatrix}\,\,,$ (A.4)
where from the second step to the third step we have assumed $P_{\epsilon_{g}\epsilon_{m}}\ll P_{\epsilon_{g}}$ and utilized the Taylor expansion $(1-x)^{-1}\sim 1+x$. For convenience, we define $X_{g}=\frac{i(\delta_{g}-b_{1}\delta)}{P_{\epsilon_{g}}}$, which is the solution of the equation of motion for $X_{g}$ when $P_{\epsilon_{g}\epsilon_{m}}=0$. Then the expression of $\boldsymbol{\phi}_{g}$ can be written as
$\bar{\boldsymbol{\phi}}_{g}=\begin{pmatrix}X_{g}+\frac{P_{\epsilon_{g}\epsilon_{m}}}{P_{\epsilon_{g}}}(2b_{1}X_{g}-\frac{i\delta}{D_{1}^{2}P_{\rm in}})\\ \frac{i\delta}{D_{1}^{2}P_{\rm in}}-b_{1}X_{g}-\frac{b_{1}P_{\epsilon_{g}\epsilon_{m}}}{P_{\epsilon_{g}}}(2b_{1}X_{g}-\frac{i\delta}{D_{1}^{2}P_{\rm in}})-\frac{P_{\epsilon_{g}\epsilon_{m}}X_{g}}{D_{1}^{2}P_{\rm in}}\\ \frac{\delta}{D_{1}}+\frac{iP_{\epsilon_{g}\epsilon_{m}}X_{g}}{D_{1}}\end{pmatrix}\,.$ (A.5)
After obtaining the expressions in Eq. (A.3), Eq. (A.5), we can substitute them into the actions Eq. (4.1) and Eq. (4.2) to calculate the contributions of the saddle points to the conditional likelihood. This has already been accomplished in [19], and we enumerate here all terms up to the third order in both $\delta$ and $\delta_{g}$.
@@
$\begin{split}(S[\bar{\boldsymbol{\phi}}]-S[\bar{\boldsymbol{\phi}}_{g}])^{(2)}={-\frac{1}{2}}\int_{\boldsymbol{k}}\frac{|{\delta_{g}(\boldsymbol{k})-\delta_{g,{\rm det}}[\delta](\boldsymbol{k})}|^{2}}{P_{\epsilon_{g}}(k)-2b_{1}P_{\epsilon_{g}\epsilon_{m}}(k)}+\Delta\wp\,\,.\end{split}$ (A.6)
@@
$\Delta\wp=\int_{\boldsymbol{k}}\frac{P_{\epsilon_{g}\epsilon_{m}}(k)\big{(}\delta_{g}(\boldsymbol{k})-\delta_{g,{\rm det}}^{(1)}[\delta](\boldsymbol{k})\big{)}\delta(-\boldsymbol{k})}{P_{\epsilon_{g}}(k)P_{\rm L}(k)}\,\,.$ (A.7)
@@
$(S[\bar{\boldsymbol{\phi}}]-S[\bar{\boldsymbol{\phi}}_{g}])^{(3)}\supset 2\int_{\boldsymbol{k}}\frac{b_{1}P_{\epsilon_{g}\epsilon_{m}}(k)}{P_{\epsilon_{g}}(k)}\frac{\big{(}\delta_{g}(\boldsymbol{k})-b_{1}\delta(\boldsymbol{k})\big{)}\,\delta^{(2)}_{g,{\rm det}}[\delta](-\boldsymbol{k})}{P_{\epsilon_{g}}(k)}\,\,,$ (A.8)

### v2

@@ 4 Example: the EFT likelihood with a Gaussian distribution
  • It has been demonstrated that the saddle-point expansion method in false vacuum decay theory can be seamlessly applied to the path-integral approach to the LSS after some appropriate modifications and re-derivations. Although we have met the negative modes, and even complex modes of $S^{\prime\prime}_{g}[\bar{\boldsymbol{\phi}}_{g}]$, because all the eigenvalues of the gradient flow eigenequation Eq. (4.14) are real and positive ($m_{n}\in\mathbb{R},m_{n}>0$), the result is ensured to be real through the path integral, consistent with the argument as presented in Sec. 2.1. This aspect of the result is very crucial, as it directly relates to its physical interpretation as a probability.
  • Although we have examined only one specific model, the rationale for applying the gradient flow equation to determine the steepest descent contours, as we discussed below, is general and can be applied to any kind of more complicated models. After this, we can compute arbitrary likelihoods of the form $\int\mathcal{D}\phi\,{\rm e}^{-S[\phi]}$ up to the second-order near the saddle points. This approach could be instrumental in providing more accurate predictions for the large-scale structure of the universe.
@@
$\wp[\delta]$ $={\rm ln}\,\mathcal{P}[\delta]=-{\rm ln}\,\left(\frac{\sqrt{\pi}}{D_{1}}\right)S[\bar{\boldsymbol{\phi}}]\,,$ (4.19)
@@
Since neither the matrices $\mathcal{M}$ nor $\mathcal{M}_{g}$ depend on the overdensity fields $\delta$ and $\delta_{g}$, the eigenvalues will be field-independent. We can also see this in the expression for $m_{n}$ in Appendix C, where we provide a set of approximate solutions for them. Thus the logarithmic terms in Eqs. (4.19) and (4.20) will be absorbed into the normalization factors. This absorption originates from the normalization factor introduced in the definition of the functional Dirac function Eq. (3.4). Consequently, this factor manifests in the expression for the likelihoods Eqs. (3.5) and (3.6).
We hereby provide a concise introduction to the renormalization scheme in this theory. Let us first consider the case in quantum field theory as an example. It appears that the application of saddle-point expansion may affect the renormalization, as the field decomposition introduces an additional fluctuation term $\varphi$, which needs a counterterm to cancel its UV dependence. However, in practice, if we aim to compute the radiative corrections to the path integral, we only need to directly use the renormalized effective action, rather than performing renormalization after the calculation of the saddle-point contributions. For example, some literature considers the 1PI [57, 36] and 2PI [39, 55] effective actions [^6] In the computation of the radiation corrections, it is often required that the effective action be convex to ensure physical consistency. A method to handle this issue using Green’s function method can be found in [39].. Returning to the present case, when there are field-independent terms in $\wp[\delta]$ and $\wp[\delta_{g},\delta]$, they can be directly ignored due to the presence of the normalization factor $\mathcal{N}_{\delta^{(\infty)}}$. When these terms exhibit field-dependence, no additional counterterm needs to be introduced when incorporating loop corrections. Instead, the renormalized “action” can be directly used.
Consequently, in this case, when defining the conditional likelihood of them, the terms appearing in the conditional likelihood will only include the saddle-point contributions of the two actions, which is
@@ 5 Summary and conclusion
The path-integral approach to the EFTofLSS offers a novel method for encapsulating cosmological effects into a compact analytical framework, enabling the prediction of cosmological observables. In this work, we aim to apply the saddle-point expansion technique, traditionally employed in quantum field theory for calculating the decay rates, to this context to improve the precision of EFT likelihood calculations. As a result, after appropriately reformulating the likelihood expressions, this task reduces to computing the path integral over the field $\boldsymbol{\phi}$ in three-dimensional Euclidean space. Consequently, by applying the same reasoning as in quantum field theory, we can calculate the conditional likelihood, with the $\phi$ space replaced by the $\boldsymbol{\phi}$ space.
@@
A key challenge in the saddle-point expansion method is the treatment of the negative eigenvalues of the second derivative of the action at the saddle point. In this work, we present a systematic discussion of this aspect in three distinct cases. As an illustrative example, we compute the matter and joint likelihood under the assumption of a Gaussian distribution, retaining terms up to $J^{2}$ and $J_{g}^{2}$ order. For the matter likelihood, since there are no negative eigenvalues associated with $S^{\prime\prime}[\boldsymbol{\phi}]$, we can directly apply the Gaussian integral formula to obtain the result. However, when calculating the joint likelihood, the eigenvalues of $S^{\prime\prime}[\boldsymbol{\phi}_{g}]$ are often non-positive and even complex, necessitating an approximation of the original integral contour with a sum of the steepest descent contours in the field space. The treatment of these complex modes goes beyond the scope of the false vacuum decay theory discussion, primarily because the second derivative of the action in the path-integral approach to the LSS often takes a specific form, resulting in a general non-Hermitian three-dimensional matrix. Even so, we still demonstrate in Sec. 4 that for this case, the Picard-Lefschetz theory remains applicable in this context by complexifying the action and using the gradient flow equation in the field space. We emphasize here again that although the integrals along each contour may yield complex results, their collective interference leads to a real final result.
@@
Another point is that, as shown in Eqs. (4.8) and (4.18), the result of the saddle-point expansion is essentially equivalent to the saddle-point approximation, differing only by a multiplicative factor. In the example provided in Sec. 4, this factor is independent of $\delta_{g}$ and $\delta$, therefore it will be absorbed into the normalization constant and resulting in the same outcome as in [19], as illustrated in Eq. (4.21). However, this does not imply that the result of the saddle-point expansion is trivial, but rather that the action we have chosen in this example is truncated at $\boldsymbol{\phi}^{2}$ order. Furthermore, the method we present in this paper is general, and can be used to cope with more complicated situations.
@@
Several unresolved issues remain in this work. For instance, we have not obtained an exact analytic solution of Eq.(4.16). The reason is that although the combination of the matrix $\mathcal{M}_{g}$ with its complex conjugate $\mathcal{M}_{g}^{*}$ is Hermitian, as illustrated in Eq. (4.16), solving for this eigenequation is equivalent to solving a cubic equation, which does not have a simple analytic solution. Another important aspect concerns the treatment of higher-order terms. In the theory of false vacuum decay, the Lagrangian is typically truncated at the $\phi^{4}$ order due to the non-renormalizable nature of higher-order terms, which are strongly suppressed by the energy scale of new physics, $\mu$. This raises the question of whether a similar suppression applies to higher-order terms in our case. In [47] and [49], the authors computed the renormalization group equations for the partition function up to $J^{2}$ order and arbitrary $J^{n}$ order ($n\rightarrow\infty$). They found that, given the same initial conditions, the running of the bias parameter $b$ exhibits similar behavior. Does this imply when calculating the EFT likelihood, we can neglect the terms higher than $J^{2}$ (analogous to the non-renormalizable terms in quantum field theory)? In [47] the authors perform a dimension analysis on the terms of all orders of $J$, however, do not draw an analogy to the case of quantum field theory. Then for the higher-order terms of $J$, are they also suppressed by the hard momentum cutoff $\Lambda$ introduced during the regularization process? e.g. consider the term $\int_{\boldsymbol{x}}J^{2}(\boldsymbol{x})O[\delta_{\Lambda}(\boldsymbol{x})]$, should we need to express this as $\frac{C_{*}}{\Lambda^{3+d_{O}}}\int_{\boldsymbol{x}}J^{2}(\boldsymbol{x})O[\delta_{\Lambda}(\boldsymbol{x})]$ [^7] Strictly speaking, this formulation is not precise. In quantum field theory, the energy scale of new physics $\mu$ and the momentum cutoff $\Lambda$ are two fundamentally distinct concepts [56]. The former is physical, introduced through dimensional regularization, and represents the energy scale at which we conduct our analysis. In contrast, the latter corresponds to the scale at which the theory breaks down in cutoff-based regularization schemes (for example, the Pauli-Villars regularization). It is often unphysical and can be taken to infinity in computations. We formulate this expression in this way to explicitly reflect the suppression of the operator. such that this term will be suppressed by $\Lambda$, where $d_{O}$ is the dimension of the arbitrary operator $O$ and $C_{*}$ is the dimensionless “coupling constant”? If similar reasoning as in quantum field theory can be applied, then these higher-order terms would contribute little to the partition function. Unfortunately, we have not found relevant discussions in any literature, so we refrain from drawing a definitive conclusion at this stage.
Additionally, our calculation assumes a Gaussian initial condition. Extending our analysis to non-Gaussian initial conditions is essential, as it will facilitate a deeper investigation of the early universe. To address the challenge that SPT fails to adequately capture primordial non-Gaussianity, one possible solution is to construct sophisticated statistical quantities [58]. Another approach is to employ functional methods similar to those presented in this work [48, 59]. In the standard approach [60, 61], if the primordial probability distribution deviates from the Gaussian form, then the effective action will include an additional interaction term. This term introduces a cubic interaction in the field $\boldsymbol{\phi}$ when calculating the likelihood. The saddle-point expansion method would then require a careful treatment of this term. We leave a more detailed investigation of this issue to future work.
@@ Acknowledgments
JYK thanks Ying-Zhuo Li for his useful discussions. We thank the anonymous reviewer for his/her professional and valuable suggestions, which help us make our work complete and rigorous. We acknowledge the support by the National Science Foundation of China (No. 12147217, 12347163), the China Postdoctoral Science Foundation (No. 2024M761110), and the Natural Science Foundation of Jilin Province, China (No. 20180101228JC).
Appendix A The explicit cancellation of the imaginary part in Eq. (2.8)
In this appendix, we present the explicit calculation of the path integral given by Eq. (2.8) and illustrate how the imaginary contributions from different $\mathcal{J}_{i}$ cancel out, ultimately yielding a real result.
In this example, what we need to calculate is
$\mathcal{Z}=\sum_{i}\mathcal{J}_{i}=\mathcal{J}_{-1}+\mathcal{J}_{0}+\mathcal{J}_{1}.$ (A.1)
After substituting the three paths, we can obtain
$\mathcal{Z}$ $=\exp\left(\frac{1}{4}\right)\int_{C_{-1}}{\rm d}z\,e^{-(z+1)^{2}}+\int_{C_{0}}{\rm d}z\,e^{\frac{1}{2}z^{2}}+\exp\left(\frac{1}{4}\right)\int_{C_{1}}{\rm d}z\,e^{-(z-1)^{2}}$
$=\exp\left(\frac{1}{4}\right)\left(\int_{-\infty}^{0}{\rm d}z\,e^{-(z+1)^{2}}+\int_{0}^{-i\infty}{\rm d}z\,e^{-(z+1)^{2}}\right)+\int_{i\infty}^{-i\infty}{\rm d}z\,e^{\frac{1}{2}z^{2}}$
$+\exp\left(\frac{1}{4}\right)\left(\int_{i\infty}^{0}{\rm d}z\,e^{-(z-1)^{2}}+\int_{0}^{\infty}{\rm d}z\,e^{-(z-1)^{2}}\right)\,,$ (A.2)
where we have already used $S^{\prime\prime}(-1)=S^{\prime\prime}(1)=1/4$ and $S^{\prime\prime}(0)=0$. We now examine the origin of the imaginary part of the integral. Firstly, the first term in the second line and the second term in the last line must be real, as both correspond to integrals along the real axis. At the same time, we can also obtain that
[…]
is real. Therefore, the only terms whose contributions are imaginary are the remaining two terms. This implies that the imaginary part of the path integral stems solely from the integration along the imaginary axis in this example. Meanwhile, it is straightforward to observe that
[…]
A comparison of the two expressions above reveals that the integrands of the two integrals are identical, while the integration intervals reversed. Consequently, the contribution from both integrals cancel each other out, which is what we expect. Therefore, the overall result of the path integral is real
$\mathcal{Z}$ $=\exp\left(\frac{1}{4}\right)\left(\int_{-\infty}^{1}{\rm d}z\,e^{-z^{2}}+\int_{-1}^{\infty}{\rm d}z\,e^{-z^{2}}\right)+\int_{-\infty}^{\infty}{\rm d}z\,e^{-\frac{1}{2}z^{2}}$
$=\exp\left(\frac{1}{4}\right)\left(\int_{-\infty}^{\infty}{\rm d}z\,e^{-z^{2}}+\int_{-1}^{1}{\rm d}z\,e^{-z^{2}}\right)+\sqrt{2\pi}$
$\approx\exp\left(\frac{1}{4}\right)\left(\sqrt{\pi}+0.75\right)+\sqrt{2\pi}\,,$ (A.6)
where, from the second line to the third line, we have substituted the approximate value of the Gaussian error function.
On the other hand, if we use the approximate formula for $\mathcal{I}_{i}$ given in Eq. (2.5) rather than the exact the integral $\mathcal{J}_{i}$, we will obtain
$\mathcal{I}_{-1}=\sqrt{\pi}\,{\rm exp}\left(\frac{1}{4}\right)\,,\,\,\,\mathcal{I}_{0}=\sqrt{-2\pi}\,,\,\,\,\mathcal{I}_{-1}=\sqrt{\pi}\,{\rm exp}\left(\frac{1}{4}\right)\,.$ (A.7)
We can observe that in this case $\mathcal{I}_{-1}$ and $\mathcal{I}_{1}$ have no imaginary part, whereas $\mathcal{I}_{0}$ exhibits the opposite behavior. Let us explain why the approximation $\mathcal{J}_{i}\sim\mathcal{I}_{i}$ fails to result in the necessary cancellation of the imaginary parts. When we approximate the original integral path by the sum of different Lefschetz thimbles, the reality and convergence of the entire integral are ensured by the Picard-Lefschetz theory [36]. However, when we replace $\mathcal{J}_{i}$ with $\mathcal{I}_{i}$, each $\mathcal{I}_{i}$ captures only the contribution from a small region around the corresponding saddle point, rather than the contribution from the entire Lefschetz thimble. This become evident from the explicit formulation of $\mathcal{I}_{i}$,
$\mathcal{I}_{i}=\sqrt{\frac{2\pi}{S^{\prime\prime}_{E}[s_{i}]}}\,e^{-S_{E}[s_{i}]}\,,$ (A.8)
which only encapsulates the local contributions in the immediate vicinity of the saddle point. This is analogous to the situation in the WKB approximation [35, 36] in quantum mechanics, where we use the classical trajectory to approximate the original one. We can check this in the context of this example: for $\mathcal{I}_{-1/1}$ associated with $s_{-1/1}$, since the integral along the path $C_{-1/1}$ around the saddle point takes along the real axis, its contribution is real, as shown in Eq. (2.9). Similarly, the integral along the path $C_{0}$ around $z=0$ is taken along the imaginary axis, hence the contribution of $\mathcal{I}_{0}$ is imaginary. We emphasize that in this work, when applying the saddle-point expansion method to evaluate the likelihood, we will use $\mathcal{J}_{i}$ rather than $I_{i}$ as our expression, as illustrated in Sec. 4.
Appendix B The contribution of the saddle-point
@@
$\frac{\delta S[\boldsymbol{\phi}]}{\delta\boldsymbol{\phi}}\bigg{|}_{\boldsymbol{\phi}=\bar{\boldsymbol{\phi}}}$ $=\mathcal{J}^{a}+\mathcal{M}^{ab}\bar{\boldsymbol{\phi}}^{b}=0\,,$ (B.1)
$\frac{\delta S_{g}[\boldsymbol{\phi}_{g}]}{\delta\boldsymbol{\phi}_{g}}\bigg{|}_{\boldsymbol{\phi}_{g}=\bar{\boldsymbol{\phi}}_{g}}$ $=\mathcal{J}^{a}_{g}+\mathcal{M}^{ab}_{g}\bar{\boldsymbol{\phi}}^{b}_{g}=0\,.$ (B.2)
These two equations can be readily solved once the forms of $\mathcal{M}$ and $\mathcal{M}_{g}$ are known, i.e. through Eqs. (4.3) and (4.4). The solution for the former case is
$\bar{\boldsymbol{\phi}}^{a}=-(\mathcal{M}^{ab})^{-1}\mathcal{J}^{b}=-\begin{pmatrix}\frac{P_{\rm in}^{-1}}{D_{1}^{2}}&\frac{1}{iD_{1}}\\ \frac{1}{iD_{1}}&0\end{pmatrix}\begin{pmatrix}i\delta\\ 0\end{pmatrix}=\begin{pmatrix}\frac{{i\delta}}{D_{1}^{2}P_{\rm in}}\\ \frac{\delta}{D_{1}}\end{pmatrix}.$ (B.3)
@@
$\bar{\boldsymbol{\phi}}_{g}^{a}$ $=-(\mathcal{M}^{ab}_{g})^{-1}\mathcal{J}^{b}_{g}=\frac{1}{-2b_{1}D_{1}^{2}P_{\epsilon_{g}\epsilon_{m}}+P_{\epsilon_{g}}D_{1}^{2}-P_{\epsilon_{g}\epsilon_{m}}P_{\rm in}^{-1}}\times$
@@
$\approx\frac{1+2b_{1}P_{\epsilon_{g}\epsilon_{m}}/P_{\epsilon_{g}}}{D_{1}^{2}P_{\epsilon_{g}}}\begin{pmatrix}iD_{1}^{2}\delta_{g}-ib_{1}D_{1}^{2}\delta-iP_{\epsilon_{g}\epsilon_{m}}P_{\rm in}^{-1}\delta\\ i\delta_{g}(-b_{1}D_{1}^{2}-P_{\epsilon_{g}\epsilon_{m}}P_{\rm in}^{-1})+i\delta(D_{1}^{2}b_{1}^{2}+P_{\epsilon_{g}}P_{\rm in}^{-1})\\ -D_{1}P_{\epsilon_{g}\epsilon_{m}}\delta_{g}+i\delta(ib_{1}D_{1}P_{\epsilon_{g}\epsilon_{m}}-iD_{1}P_{\epsilon_{g}})\end{pmatrix}\,\,,$ (B.4)
where from the second step to the third step we have assumed $P_{\epsilon_{g}\epsilon_{m}}\ll P_{\epsilon_{g}}$ and used the Taylor expansion $(1-x)^{-1}\sim 1+x$. For convenience, we define $X_{g}=\frac{i(\delta_{g}-b_{1}\delta)}{P_{\epsilon_{g}}}$, which is the solution of the equation of motion for $X_{g}$ when $P_{\epsilon_{g}\epsilon_{m}}=0$. Then the expression for $\boldsymbol{\phi}_{g}$ can be written as
$\bar{\boldsymbol{\phi}}_{g}=\begin{pmatrix}X_{g}+\frac{P_{\epsilon_{g}\epsilon_{m}}}{P_{\epsilon_{g}}}(2b_{1}X_{g}-\frac{i\delta}{D_{1}^{2}P_{\rm in}})\\ \frac{i\delta}{D_{1}^{2}P_{\rm in}}-b_{1}X_{g}-\frac{b_{1}P_{\epsilon_{g}\epsilon_{m}}}{P_{\epsilon_{g}}}(2b_{1}X_{g}-\frac{i\delta}{D_{1}^{2}P_{\rm in}})-\frac{P_{\epsilon_{g}\epsilon_{m}}X_{g}}{D_{1}^{2}P_{\rm in}}\\ \frac{\delta}{D_{1}}+\frac{iP_{\epsilon_{g}\epsilon_{m}}X_{g}}{D_{1}}\end{pmatrix}\,.$ (B.5)
After obtaining the expressions in Eqs. (B.3) and (B.5), we can substitute them into the actions Eqs. (4.1) and (4.2) to calculate the contributions of the saddle points to the conditional likelihood. This has already been accomplished in [19], and we enumerate here all terms up to the third order in both $\delta$ and $\delta_{g}$.
@@
$\begin{split}(S[\bar{\boldsymbol{\phi}}]-S[\bar{\boldsymbol{\phi}}_{g}])^{(2)}={-\frac{1}{2}}\int_{\boldsymbol{k}}\frac{|{\delta_{g}(\boldsymbol{k})-\delta_{g,{\rm det}}[\delta](\boldsymbol{k})}|^{2}}{P_{\epsilon_{g}}(k)-2b_{1}P_{\epsilon_{g}\epsilon_{m}}(k)}+\Delta\wp\,\,.\end{split}$ (B.6)
@@
$\Delta\wp=\int_{\boldsymbol{k}}\frac{P_{\epsilon_{g}\epsilon_{m}}(k)\big{(}\delta_{g}(\boldsymbol{k})-\delta_{g,{\rm det}}^{(1)}[\delta](\boldsymbol{k})\big{)}\delta(-\boldsymbol{k})}{P_{\epsilon_{g}}(k)P_{\rm L}(k)}\,\,.$ (B.7)
@@
$(S[\bar{\boldsymbol{\phi}}]-S[\bar{\boldsymbol{\phi}}_{g}])^{(3)}\supset 2\int_{\boldsymbol{k}}\frac{b_{1}P_{\epsilon_{g}\epsilon_{m}}(k)}{P_{\epsilon_{g}}(k)}\frac{\big{(}\delta_{g}(\boldsymbol{k})-b_{1}\delta(\boldsymbol{k})\big{)}\,\delta^{(2)}_{g,{\rm det}}[\delta](-\boldsymbol{k})}{P_{\epsilon_{g}}(k)}\,\,,$ (B.8)